# Beyond the random phase approximation for calculating Curie temperatures in ferromagnets: application to Fe, Ni, Co and monolayer CrI3

### v1

@@ III.2.2 Critical Temperatures
The Curie temperature of bcc Fe, fcc Co and fcc Ni are obtained from MFA, RPA, CD and HP are tabulated in Tab. 1. As these compounds are metallic, the spin in site-based models are best regarded as continuous variables and the calculations were thus performed in the classical limit. In order to due this for the HP and Callen methods, we calculate $T_{\mathrm{c}}(S)/S^{2}$ in the limit of $S\rightarrow\infty$ and then obtain the critical value by multiplying by the actual site-spin obtained from first principles. In the case of RPA, the classical limit can be obtained from Eq. (24) by replacing $S(S+1)$ by $S^{2}$. We state the numbers obtained from bare MFT (in brackets) as well as renormalized MFT. Comparing the TDDFT calculations of these materials [39] shows that the magnetic excitation spectrum of fcc Co is by far the simplest in the sense that the spectral function is well approximated by a Lorentzian line-shape centered on a valued, which is in excellent agreement with the prediction from renormalized MFT. We thus regard the exchange parameters and magnon dispersion of fcc Co as the most reliable. It is observed that the Curie temperature of fcc Co is in agreement with the experimental value if one uses MFA combined with bare MFT. This is a good example of error cancellation that could lead one to believe that the combination of these two approaches provide a good account of the magnetic properties. However, from Fig. 1 it is clear that MFA is expected to overestimate the Curie temperature rather severely. Indeed, when applying the more accurate renormalized MFT dispersion, MFA leads to an overestimation of the critical temperature by 28 %, which is significant although less severe than for the the simple cubic lattice due to the higher coordination number in the fcc lattice. On the other hand, the RPA, CD and HP methods predict Curie temperatures rather close to the experimental one when based on the renormalized MFT, whereas they underestimate it when combined with bare MFT (due to underestimated magnon energies).
@@
The situation is somewhat more complicated for Fe and Ni, where the renormalized MFT results do not provide as accurate an account of the magnetic excitations (compared to TDDFT) in the entire Brillouin zone. For Fe, the MFA, based on renormalized MFT, overestimates the Curie temperature whereas RPA, HP and CD underestimates it. The failure of the latter methods is likely related to fact that the transverse susceptibility of Fe extends to high energies (up to 1 eV) in large regions of the Brillouin zone (not captured by MFT) due to significant mode branching and deviation form Lorentzian lineshapes in these regions. Again, we observe that the error is smaller in the HP and CD methods compared to RPA, in agreement with the cubic model of Sec. III.1. The closest agreement with the experimental value is obtained with the MFA based on bare MFT energies, but again relies on error cancellation between underestimated exchange parameters and overestimated Curie temperatures in the MFA. For fcc Ni the situation is similar except that the HP method seems to have the worst performance. The fact that DFT based methods (combined with RPA) underestimate the Curie temperature has been discussed at some length in the literature [22, 23] and has been used to argue for the importance of renormalization. However, the experimental dispersion relation for Ni appears to be in better agreement with bare MFT [23]. Nevertheless, it is clear that the renormalized MFT dispersion is much closer to results obtained from TDDFT (in the adiabatic local density approximation) [39] and the overestimated magnon dispersion in renormalized MFT is likely because LDA itself fails to provide an accurate representation of the ground state [40, 41].
@@
We now turn to 2D systems where magnetic anisotropy plays a crucial role for magnetic order at finite temperature and an accurate decoupling of the single-ion anisotropy therefore becomes decisive. In Ref. [35] it was shown that the CD scheme compares well with an exact treatment if the anisotropy constant is not too large relative to the exchange constants, which is typically satisfied in real materials. On the other hand, it was also shown that for low spin the CD performs poorly compared to RPA in isotropic systems, leading to a suggestion of mixing the schemes (RPA+CD), such that the all terms except the single-ion anisotropy are decoupled according to (II.3) and the corrections from single-ion anisotropy are included through the Callen decoupling Eq. (II.5).
@@
In Fig. 3 we show the critical temperatures obtained in the classical limit for hexagonal, square and honeycomb lattices with nearest neighbor interaction $J$. We refer to Appendix A for a generalization of the methods to an arbitrary number of magnetic sites in the unit cell (we need this for the two sites of the honeycomb lattice). We display the results calculated with HP, RPA, RPA+CD as well as the classical MC results, which may be regarded as the exact reference. In all cases we observe the HP method to underestimate critical temperature although it acquires the best agreement with MC when $A/J$ becomes sufficiently large. In addition, RPA yields the largest deviation for large anisotropies, while the CD approach yields the worst results for small anisotropies. The RPA+CD approach performs better than both the RPA and CD for any value of the anisotropy, but performs significantly worse than the HP method for large anisotropies. It has previously been shown that all the decoupling schemes fail rather dramatically in the limit of $A/J\rightarrow\infty$. In that case the transverse fluctuations are frozen out and the Heisenberg Hamiltonian reduces to the Ising model, but all of the decoupling schemes yields $T_{\mathrm{c}}\propto A$ for $A/J\rightarrow\infty$ [9, 35] rather than $T_{\mathrm{c}}\rightarrow T_{\mathrm{c}}^{\mathrm{Ising}}$. The origin of this behaviour can be exemplified by the MFA where a the mean-field decoupling of the single-ion anisotropy yields $3k_{\mathrm{B}}T_{\mathrm{c}}=S(S+1)(J_{0}+2A)$, while the exact treatment must become independent of $A$ in the limit of $A/J_{0}\rightarrow\infty$. Here we just note that the Curie temperatures obtained from RPA exhibit the fastest increase with respect to $A$ and breaks the Ising limit rather quickly, whereas the HP method has the slowest increase in Curie temperature and only breaks the Ising limit at rather a rather large value of $A/J$. Although we have only compared with exact results in the classical limit, these calculations could indicate that the RPA+CD method has a superior performance for 2D systems with single-ion anisotropy in the regime where $A/J<0.2$.
@@ III.4 Application to monolayer CrI3
Figure 4: Curie temperatures of CrI3 obtained with nearest neighbor interactions $J_{1}=2.0$ meV. The results are shown as a function of $J_{2}$ for $\lambda^{12}_{0}=0$ and $A=0$ with the remaining anisotropy parameters obtained from the experimental spinwave gap.
Here we will not delve into different methodologies for obtaining exchange constants but rather focus on the different approaches for calculating the Curie temperature given a set of Heisenberg parameters. Although these are not known exactly, the spinwave gap and the optical magnon at the $\Gamma$-point for bilayer CrI3 has been measured to be 0.3 meV and 18.4 meV respectively and these values puts strong bounds on the exchange constants. In particular, within the model defined by Eqs. (1) and (32), the spinwave gap at $T=0$ obtained in either semi-classical spinwave theory (relevant for classical MC) or RPA is given by
@@
where $J^{12}_{0}$ and $\lambda^{12}_{0}$ are the sum of inter-sublattice exchange constants and anisotropic exchange constants respectively (defined in Eqs.(56) and (57)). We have ignored the effect of of intra-sublattice anisotropic exchange($\lambda^{11}_{0}$) in our model. Most first principles methods agree that the exchange constants vanish beyond third nearest neighbors and the isotropic model is thus defined by $J_{1}$, $J_{2}$ and $J_{3}$, which yields $J^{12}_{0}=N_{\mathrm{nn}}(J_{1}+J_{3})$ with $N_{\mathrm{nn}}=3$. The third nearest neighbor exchange is typically predicted to be rather small and we may thus use Eq. (48) to get $J_{1}=2.0$ meV. Similarly, depending on whether one uses RPA/MC or HP/CD for modelling, the single-ion anisotropy and sum of anisotropic exchange constants may be obtained from Eq. (46) or Eq. (47) respectively. Now, it is not clear whether the spin-wave gap in CrI3 originates from either of these constants. Specifically, LDA and PBE predict equal contribution from $A$ and $\lambda^{12}_{0}$ while LDA+U and PBE+U predict that $A\sim 0$ and the spin-wave gap thus mainly originate from anisotropic exchange. However, all of these functionals predict a spin-wave gap of $\sim 0.7-1.0$ meV, which is more than a factor of two larger than the experimental gap and the reliability of any of these predictions is thus questionable.
@@
In Fig. 4 we show the critical temperatures obtained with classical MC, HP, RPA, CD and RPA+CD as a function of $J_{2}$ using the experimental value of $J_{1}$ discussed above. We show the results for the two limiting cases of $\lambda^{12}_{0}=0$ and $A=0$ where the remaining anisotropy parameter is obtained from Eqs. (46) or (47). The anisotropic exchange in these models are limited to nearest neighbors. We first note that the Curie temperature in all methods except for the HP is largely independent on whether the spin wave gap originates from single-ion anisotropy or anisotropic exchange. We believe that this represents a severe problem with the description of single-ion anisotropy in the HP method. In particular, one may show that the Curie temperature in the HP method will decrease with increasing anisotropic exchange (and increasing spin wave gap) for a fixed value of $A$, which seems unreasonable. The thermal renormalization of the spin-wave gap is included through the term $-4A\phi$ in Eq. (34), which leads to a (too) rapid closing of the gap at elevated temperatures. In contrast, the renormalization is rescaled by $\langle S^{z}\rangle^{2}$ in the CD method, which leads to a slower and continuous closing of the gap when the temperature is increased.
@@
Based on the classical limit illustrated in Fig. 3 as well as previous simulations on spin-1/2 systems, it appears likely that the RPA+CD method comprises the most accurate approach. Based on the present results it is, however, not possible to assert this statement, since the exact value of $J_{2}$ is not known. LDA calculations yield $J_{1}-J_{3}=2.0$ meV and $J_{2}=0.6$ meV, while more accurate calculations based on many-body perturbation theory [42] as well as experiments on bulk CrI3 [43, 44] imply that $J_{2}$ should be on the order of 0–0.2 meV. However, the magnetic excitation spectrum contains a significant gap at the $K$-point as well [43, 44], which cannot be accounted for by the present model [42, 45]. Nonetheless, we wish to emphasize the uncertainty in obtaining critical temperatures for 2D materials from simple decoupling schemes such as the ones studied here, where the predicted Curie temperature can differ by more than a factor of two. It is therefore highly unreliable to benchmark first principles simulations against experimental Curie temperatures of 2D materials obtained by the methods discussed here. The worst approach for non-classical systems is arguably the classical MC, which severely underestimates the Curie temperature. This can be understood from the MFA in 3D, which will predict 60 % of the quantum value for a spin-3/2 system and it is highly likely that such errors are inherited by a classical treatment of 2D systems. In addition, the HP method appears to become unreliable when single-ion anisotropy is included.
@@ IV Conclusion and outlook
We have compared the performance of classical MC, MFA, RPA, HP the CD decoupling schemes for calculating Curie temperatures of ferromagnets. In the classical limit, the MC simulations are regarded as the exact reference and for a cubic lattice with nearest neighbor isotropic interactions, the CD method becomes highly accurate while the HP overestimates the Curie temperatures slightly and RPA underestimates it. In contrast, the MFA yields a significant overestimation (39 %) of the Curie temperatures in the classical limit. It has, however, previously been shown that the RPA method is more accurate than the CD for $S=1/2$ systems [35]. On the other hand, from Fig. 1 it is clear that the HP decoupling interpolates between RPA and the CD results when $S$ is increased from $1/2$ to the classical limit and the HP method could perhaps be regarded as the best comprise for isotropic systems with arbitrary spin.
@@
Such an approach would take the lineshape of the magnon modes as well as possible mode branching into account, but would comprise a tremendous computational task due to the requirement of densely sampled TDDFT calculations in the entire Brillouin zone. Nevertheless, it would definitely be highly enlightening to pursue.
@@
For the case of 2D materials, the conclusions become even more blurry. The reason is that magnetic anisotropy plays a crucial role for magnetic order and the inclusion of such terms in the Heisenberg model makes the decoupling schemes more dubious. By default, RPA fails to reproduce the lack of magnetic order for $S=1/2$ systems with single-ion anisotropy, while the HP method appears to severely overestimate the thermal renormalization of the spin-wave gap, whereas the CD method overestimates Curie temperatures when the spin-wave gap becomes small. We have exemplified these problems in the classical limit for honeycomb, square and hexagonal lattices where we may again compare to classical MC calculations. In all cases, the RPA+CD (RPA method for isotropic and CD for anisotropic terms) yields good agreement with classical MC for low values of anisotropy ($A/J<0.2$). For large anisotropies, all the methods break down and asymptotically approach a linear relation between the single-ion anisotropy and the critical temperature, whereas the exact result should approach the Ising limit asymptotically. In the quantum regime, on the other hand, the classical MC leads to significant underestimation of critical temperature and is not a viable alternative to the quantum mechanical decoupling schemes.
@@
We applied the methods for obtaining Curie temperatures in 2D to the case of CrI3, where we focused on Heisenberg parameters extracted from experiments. This constitutes a rather incomplete description of the system and we left the next nearest exchange coupling as a free parameter. It is, however, evident that the predicted Curie temperatures are rather sensitive to the applied method. It is difficult to judge which of the Green’s function methods (RPA, CD and RPA+CD) that is most appropriate for CrI3, but based on the previous calculations we believe that RPA+CD should be the preferred approach. The HP method, yields significantly lower Curie temperatures (comparable to classical MC) if order is assumed driven by single-ion anisotropy, whereas it seems to provide the best agreement with experiments if order is driven by anisotropic exchange. We believe that the HP probably constitutes the most accurate method if single-ion anisotropy can be neglected, but it should be avoided if single-ion anisotropy plays a prominent role. For classical MC, the Curie temperature is expected to be underestimated in general. As a result, previous reports on theoretical critical temperature that are in agreement with experimental ones, are likely prone to a fortuitous error cancellation (due to inaccurate overestimated exchange constants) whenever these are based on classical MC combined with DFT [9, 4, 47, 48].
@@
In this way, the accuracy of different decoupling schemes in 3D is not directly transferable to 2D magnets due to the inherent problems associated with decoupling the single-ion anisotropy from the spin-flip propagation. While there are plenty of bulk 3D materials where Heisenberg parameters have been determined rigorously from inelastic neutron scattering, the parameters for 2D magnets are more scarce. The acquisition of such data is crucial for properly benchmarking the performance of different decoupling schemes in 2D and the experimental extraction of accurate parameters thus seems pertinent for computational progress in the field. Nevertheless, based on the results in the present paper as well as previous works we believe that one may formulate recommendations that depend on the type of system under scrutiny. In Tab. 2, we summarize the methodologies that are expected to be most accurate for different types of materials and hope that this may serve as useful guidelines for future calculations.

### v2

@@ III.2.2 Critical Temperatures
The Curie temperature of bcc Fe, fcc Co and fcc Ni obtained from MFA, RPA, CD and HP are tabulated in Tab. 1. As these compounds are metallic, the spin in site-based models are best regarded as continuous variables and the calculations were thus performed in the classical limit. In order to due this for the HP and Callen methods, we calculate $T_{\mathrm{c}}(S)/S^{2}$ in the limit of $S\rightarrow\infty$ and then obtain the critical value by multiplying by the actual site-spin obtained from first principles. In the case of RPA, the classical limit can be obtained from Eq. (24) by replacing $S(S+1)$ by $S^{2}$. We state the numbers obtained from bare MFT (in brackets) as well as renormalized MFT. Comparing the TDDFT calculations of these materials [39] shows that the magnetic excitation spectrum of fcc Co is by far the simplest in the sense that the spectral function is well approximated by a Lorentzian line-shape centered on a value, which is in excellent agreement with the prediction from renormalized MFT. We thus regard the exchange parameters and magnon dispersion of fcc Co as the most reliable. It is observed that the Curie temperature of fcc Co is in agreement with the experimental value if one uses MFA combined with bare MFT. This is a good example of error cancellation that could lead one to believe that the combination of these two approaches provide a good account of the magnetic properties. However, from Fig. 1 it is clear that MFA is expected to overestimate the Curie temperature rather severely. Indeed, when applying the more accurate renormalized MFT dispersion, MFA leads to an overestimation of the critical temperature by 28 %, which is significant although less severe than for the the simple cubic lattice due to the higher coordination number in the fcc lattice. On the other hand, the RPA, CD and HP methods predict Curie temperatures rather close to the experimental one when based on the renormalized MFT, whereas they underestimate it when combined with bare MFT (due to underestimated magnon energies).
@@
The situation is somewhat more complicated for Fe and Ni, where the renormalized MFT results do not provide as accurate an account of the magnetic excitations (compared to TDDFT) in the entire Brillouin zone. For Fe, the MFA, based on renormalized MFT, overestimates the Curie temperature whereas RPA, HP and CD underestimates it. The failure of the latter methods is likely related to fact that the transverse susceptibility of Fe extends to high energies (up to 1 eV) in large regions of the Brillouin zone (not captured by MFT) due to significant mode branching and deviation from Lorentzian lineshapes in these regions. Again, we observe that the error is smaller in the HP and CD methods compared to RPA, in agreement with the cubic model of Sec. III.1. The closest agreement with the experimental value is obtained with the MFA based on bare MFT energies, but again relies on error cancellation between underestimated exchange parameters and overestimated Curie temperatures in the MFA. For fcc Ni the situation is similar except that the HP method seems to have the worst performance. The fact that DFT based methods (combined with RPA) underestimate the Curie temperature has been discussed at some length in the literature [22, 23] and has been used to argue for the importance of renormalization. However, the experimental dispersion relation for Ni appears to be in better agreement with bare MFT [23]. Nevertheless, it is clear that the renormalized MFT dispersion is much closer to results obtained from TDDFT (in the adiabatic local density approximation) [39] and the overestimated magnon dispersion in renormalized MFT is likely because LDA itself fails to provide an accurate representation of the ground state [40, 41].
@@
We now turn to 2D systems where magnetic anisotropy plays a crucial role for magnetic order at finite temperature and an accurate decoupling of the single-ion anisotropy therefore becomes decisive. In Ref. [35] it was shown that the CD scheme compares well with an exact treatment if the anisotropy constant is not too large relative to the exchange constants, which is typically satisfied in real materials. On the other hand, it was also shown that for low spin the CD performs poorly compared to RPA in isotropic systems, leading to a suggestion of mixing the schemes (RPA+CD), such that all terms except the single-ion anisotropy are decoupled according to (II.3) and the corrections from single-ion anisotropy are included through the Callen decoupling Eq. (II.5).
@@
In Fig. 3 we show the critical temperatures obtained in the classical limit for hexagonal, square and honeycomb lattices with nearest neighbor interaction $J$. We refer to Appendix A for a generalization of the methods to an arbitrary number of magnetic sites in the unit cell (we need this for the two sites of the honeycomb lattice). We display the results calculated with HP, RPA, RPA+CD as well as the classical MC results. Even though classical MC may serve as the exact reference, simulations in the limit of $A/J\rightarrow 0$e difficult to converge and less reliable because the $T_{\mathrm{c}}$ predicted by heat capacity and magnetization starts to diverge. So, we limited ourselves to a range where classical MC is reliable. In all cases we observe the HP method to underestimate critical temperature although it acquires the best agreement with MC when $A/J$ becomes sufficiently large. In addition, RPA yields the largest deviation for large anisotropies, while the CD approach yields the worst results for small anisotropies. The RPA+CD approach performs better than both the RPA and CD for any value of the anisotropy, but performs significantly worse than the HP method for large anisotropies. It has previously been shown that all the decoupling schemes fail rather dramatically in the limit of $A/J\rightarrow\infty$. In that case the transverse fluctuations are frozen out and the Heisenberg Hamiltonian reduces to the Ising model, but all of the decoupling schemes yields $T_{\mathrm{c}}\propto A$ for $A/J\rightarrow\infty$ [9, 35] rather than $T_{\mathrm{c}}\rightarrow T_{\mathrm{c}}^{\mathrm{Ising}}$. The origin of this behaviour can be exemplified by the MFA where the mean-field decoupling of the single-ion anisotropy yields $3k_{\mathrm{B}}T_{\mathrm{c}}=S(S+1)(J_{0}+2A)$, while the exact treatment must become independent of $A$ in the limit of $A/J_{0}\rightarrow\infty$. Here we just note that the Curie temperatures obtained from RPA exhibit the fastest increase with respect to $A$ and breaks the Ising limit rather quickly, whereas the HP method has the slowest increase in Curie temperature and only breaks the Ising limit at a rather rather large value of $A/J$. Although we have only compared with exact results in the classical limit, these calculations could indicate that the RPA+CD method has a superior performance for 2D systems with single-ion anisotropy in the regime where $A/J<0.2$.
@@ III.4 Application to monolayer CrI3
Figure 4: Curie temperatures of CrI3 obtained with nearest neighbor interactions $J_{1}=2.0$ meV and anisotropy parameters consistent with the experimental spin wave gap of $\Delta=0.3$ meV. Left: Curie temperature as function of $J_{2}$ with $\lambda^{12}_{0}=0$ and $A$ determined from Eqs. (46)-(47). Center: Curie temperature as function of $J_{2}$ with $A=0$ and $\lambda^{12}_{0}$ determined from Eqs. (46)-(47). Right: Curie temperature as a function of the single-ion anisotropy $A$ with $J_{2}=0$ and $\lambda^{12}_{0}$ determined from Eq. (47).
We will not delve into different methodologies for obtaining exchange constants but only focus on the different approaches for calculating the Curie temperature given a set of Heisenberg parameters. Although these are not known exactly, the spinwave gap and the optical magnon at the $\Gamma$-point for bilayer CrI3 have been determined experimentally to be 0.3 meV and 18.4 meV respectively [42] and these values puts strong bounds on the exchange constants. In particular, within the model defined by Eqs. (1) and (32), the spinwave gap at $T=0$ obtained in either semi-classical spinwave theory (relevant for classical MC) or RPA is given by
@@
where $J^{12}_{0}$ and $\lambda^{12}_{0}$ are the sum of inter-sublattice exchange constants and anisotropic exchange constants respectively (defined in Eqs.(56) and (57)). We have ignored the effect of of intra-sublattice anisotropic exchange ($\lambda^{11}_{0}$) in our model. Most first principles methods agree that the exchange constants vanish beyond third nearest neighbors and the isotropic model is thus defined by $J_{1}$, $J_{2}$ and $J_{3}$, which yields $J^{12}_{0}=N_{\mathrm{nn}}(J_{1}+J_{3})$ with $N_{\mathrm{nn}}=3$. The third nearest neighbor exchange is typically predicted to be rather small and we may thus use Eq. (48) to get $J_{1}=2.0$ meV, while $J_{2}$ remains undetermined by experiements. Similarly, depending on whether one uses RPA/MC or HP/CD for modelling, the single-ion anisotropy and sum of anisotropic exchange constants may be obtained from Eq. (46) or Eq. (47) respectively. Now, it is not clear whether the spin-wave gap in CrI3 originates from either of these constants. Specifically, LDA and PBE predict equal contribution from $A$ and $\lambda^{12}_{0}$ while LDA+U and PBE+U predict that $A\sim 0$ and the spin-wave gap thus mainly originate from anisotropic exchange [30]. However, all of these functionals predict a spin-wave gap of $\sim 0.7-1.0$ meV, which is more than a factor of two larger than the experimental gap and the reliability of any of these predictions may be questionable.
@@
In Fig. 4 we show the dependence of the critical temperatures obtained with classical MC, HP, RPA, CD and RPA+CD, while the parameters $J_{2}$, $\lambda^{12}_{0}$ and $A$ are varied. In all cases we have taken $J_{1}=2.0$ meV, $J_{3}=0$ (consistent with the optical magnon at $\Gamma$) and with the spin wave gap fixed to the experimental value of 0.3 meV. In the first two plots, we show the results for the two limiting cases of $\lambda^{12}_{0}=0$ and $A=0$ where the remaining anisotropy parameter is obtained from Eqs. (46) or (47). We also show the variation when the anisotropy (represented by the experimental spin wave gap) is mapped to different values of $A$(and $\lambda^{12}_{0})$, while enforcing $\Delta=0.3$ meV and $J_{2}=0$. We first note that the Curie temperature in the classical MC, RPA (not shown, but invariant by construction) and CD approaches are largely independent on whether the spin wave gap originates from single-ion anisotropy or anisotropic exchange. The RPA+CD exhibits moderate variation, while the HP method has strong dependence on the origin of anisotropy. We believe that this represents a severe problem with the description of single-ion anisotropy in the HP method. In particular, one may show that the Curie temperature in the HP method will decrease with increasing anisotropic exchange (and increasing spin wave gap) for a fixed value of $A$, which seems unreasonable. The thermal renormalization of the spin-wave gap is included through the term $-4A\phi$ in Eq. (34), which leads to a (too) rapid closing of the gap at elevated temperatures. In contrast, the renormalization is rescaled by $\langle S^{z}\rangle^{2}$ in the CD method, which leads to a slower and continuous closing of the spinwave gap when the temperature is increased.
@@
Based on the classical limit illustrated in Fig. 3 as well as previous simulations on spin-1/2 systems, it appears likely that the RPA+CD method comprises the most accurate approach. Based on the present results it is, however, not possible to assert this statement, since the exact value of $J_{2}$ is not known. LDA calculations yield $J_{1}+J_{3}=2.0$ meV and $J_{2}=0.6$ meV, while more accurate calculations based on many-body perturbation theory [43] as well as experiments on bulk CrI3 [44, 45] imply that $J_{2}$ should be on the order of 0–0.2 meV. However, the magnetic excitation spectrum contains a significant gap at the $K$-point as well [44, 45], which cannot be accounted for by the present model [43, 46]. Nonetheless, we wish to emphasize the uncertainty in obtaining critical temperatures for 2D materials from simple decoupling schemes such as the ones studied here, where the predicted Curie temperature can differ by more than a factor of two depending on the method of choice. It may therefore be unreliable to benchmark first principles simulations against experimental Curie temperatures of 2D materials obtained by the methods discussed here. The worst approach for non-classical systems is arguably the classical MC, which severely underestimates the Curie temperature. This can be understood from the MFA in 3D, which will predict 60 % of the quantum value for a spin-3/2 system and it is highly likely that such errors are inherited by a classical treatment of 2D systems. In addition, the HP method appears to become unreliable when single-ion anisotropy is included.
@@ IV Conclusion and outlook
We have compared the performance of classical MC, MFA, RPA, HP and the CD decoupling schemes for calculating Curie temperatures of ferromagnets. In the classical limit, the MC simulations are regarded as the exact reference and for a cubic lattice with nearest neighbor isotropic interactions, the CD method becomes highly accurate while the HP overestimates the Curie temperatures slightly and RPA underestimates it. In contrast, the MFA yields a significant overestimation (39 %) of the Curie temperatures in the classical limit. It has, however, previously been shown that the RPA method is more accurate than the CD for $S=1/2$ systems [35]. On the other hand, from Fig. 1 it is clear that the HP decoupling interpolates between RPA and the CD results when $S$ is increased from $1/2$ to the classical limit and the HP method could perhaps be regarded as the best comprise for isotropic systems with arbitrary spin.
@@
Such an approach would take the lineshape of the magnon modes as well as possible mode branching into account, but would comprise a tremendous computational task due to the requirement of densely sampled TDDFT calculations in the entire Brillouin zone.
@@
For the case of 2D materials, the conclusions become a bit more more blurry. The reason is that magnetic anisotropy plays a crucial role for magnetic order and the inclusion of such terms in the Heisenberg model makes the decoupling schemes more dubious. By default, RPA fails to reproduce the lack of magnetic order for $S=1/2$ systems with single-ion anisotropy, while the HP method appears to severely overestimate the thermal renormalization of the spin-wave gap and the CD method overestimates Curie temperatures when the spin-wave gap becomes small. We have exemplified these problems in the classical limit for honeycomb, square and hexagonal lattices where we again compare to classical MC calculations. In all cases, the RPA+CD (RPA method for isotropic and CD for anisotropic terms) yields good agreement with classical MC for low values of anisotropy ($A/J<0.2$). For large anisotropies, all the methods break down and asymptotically approach a linear relation between the single-ion anisotropy and the critical temperature, whereas the exact result should approach the Ising limit asymptotically. On the other hand, in the quantum regime the classical MC leads to significant underestimation of critical temperature and therefore does not comprise a viable alternative to the quantum mechanical decoupling schemes.
@@
We applied the methods for obtaining Curie temperatures in 2D to the case of CrI3, where we focused on Heisenberg parameters extracted from experiments. This constitutes a rather incomplete description of the system and we left the next nearest exchange coupling as a free parameter. It is, however, evident that the predicted Curie temperatures are rather sensitive to the applied method. It is difficult to judge which of the Green’s function methods (RPA, CD and RPA+CD) that is most appropriate for CrI3, but based on the previous calculations we believe that RPA+CD should be the preferred approach. The HP method, yields significantly lower Curie temperatures (comparable to classical MC) if the magnetic anisotropy is governed by single-ion anisotropy, whereas it seems to provide the best agreement with experiments if the anisotropy is governed by anisotropic exchange. We believe that the HP probably constitutes the most accurate method if single-ion anisotropy can be neglected, but it should be avoided if single-ion anisotropy plays a prominent role. For classical MC, the Curie temperature is expected to be underestimated in general. As a result, previous reports on theoretical critical temperature that are in agreement with experimental ones, are likely prone to a fortuitous error cancellation (due to overestimated exchange constants) whenever these are based on classical MC combined with DFT [9, 4, 48, 49].
@@
In this way, the accuracy of different decoupling schemes in 3D is not directly transferable to 2D magnets due to the inherent problems associated with decoupling the single-ion anisotropy. While there are plenty of bulk 3D materials where Heisenberg parameters have been determined rigorously from inelastic neutron scattering, the parameters for 2D magnets are more scarce. The acquisition of such data is crucial for properly benchmarking the performance of different decoupling schemes in 2D and the experimental extraction of accurate parameters thus seems pertinent for computational progress in the field. Nevertheless, based on the results in the present paper as well as previous works we believe that one may formulate recommendations that depend on the type of system under scrutiny. In Tab. 2, we summarize the methodologies that are expected to be most accurate for different types of materials and hope that this may serve as useful guidelines for future calculations.